RENNER-TELLER EFFECTS IN THE ^{2}Π_{i} GROUND STATES OF THE SiCH/SiCD and GeCH/GeCD RAD-ICALS

\underline{HAIYANG LI}, TONY C. SMITH, DAVID A. HOSTUTLER, DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055; ANTHONY J. MERER, Chemistry Department, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1.

Extensive emission spectra have been obtained by laser excitation of selected bands of the Ã ^{2}Σ^{+} - X̃ ^{2}Π_{i} systems of jet-cooled SiCH, SiCD, GeCH and GeCD. Up to four quanta of the Si-C or Ge-C stretching mode and even quanta of the bending mode have been observed by pumping upper state Σ vibronic levels. Pumping vibronically induced transitions to upper state Π vibronic levels has given further information about odd quanta of the bending mode. The assignments have been fitted using a vibronic coupling matrix that includes spin-orbit coupling, Renner-Teller effects, Fermi resonances, and the interaction between nearly degenerate bending levels of different ν_{2} and K. The emission spectra of SiCH and SiCD, with A ≈ ϵω, are easily assigned and the Fermi resonance interaction between 2ν_{2} and ν_{3} is found to be small. In the germanium species, the spin-orbit coupling is much larger than ϵω and the spectra are more complicated. Levels of different ν_{2} with Δν_{2} = ±1 are mixed by a ΔP = 0, H_{RT} ×H_{SO} cross-term of substantial magnitude, disrupting the energy level pattern. Results from the vibronic analyses of all four species will be presented.